OpenStudy (august899):

Complete the following identity. ________?=sec^2 θ (Multiple Choice)

OpenStudy (august899):

A. tan^2 θ+1 B.tan^2θ-1 C.1+cot^2 θ D.1-cot^2 θ

OpenStudy (ifrah123):

A

OpenStudy (august899):

Can you show me how to work that out?

jimthompson5910 (jim_thompson5910):

Start with the pythagorean identity (step 1). Then divide both sides by cos^2(theta) (step 2). On step 3, you'll break up the fraction. Then on step 4, you use these identities tan = sin/cos sec = 1/cos x/x = 1 where x is anything but zero So doing all that means we have these steps... \[\Large \sin^2(\theta)+\cos^2(\theta) = 1\] \[\Large \frac{\sin^2(\theta)+\cos^2(\theta)}{\cos^2(\theta)} = \frac{1}{\cos^2(\theta)}\] \[\Large \frac{\sin^2(\theta)}{\cos^2(\theta)}+\frac{\cos^2(\theta)}{\cos^2(\theta)} = \frac{1}{\cos^2(\theta)}\] \[\Large \tan^2(\theta)+1 = \sec^2(\theta)\]

OpenStudy (marcelie):

you can also work those problems out but you can memorize them too